# Improving Handwritten Text Recognition via 3D Attention and Multi-Scale Training

### v2

@@ Abstract
The segmentation-free research efforts for addressing handwritten text recognition can be divided into three categories: connectionist temporal classification (CTC), hidden Markov model and encoder–decoder methods. In this paper, inspired by the above three modeling methods, we propose a new recognition network by using a novel three-dimensional (3D) attention module and global-local context information. Based on the feature maps of the last convolutional layer, a series of 3D blocks with different resolutions are split. Then, these 3D blocks are fed into the 3D attention module to generate sequential visual features. Finally, by integrating the visual features and the corresponding global-local context features, a well-designed representation can be obtained. Main canonical neural units including attention mechanisms, fully-connected layer, recurrent unit and convolutional layer are efficiently organized into a network and can be jointly trained by the CTC loss and the cross-entropy loss. Experiments on the latest Chinese handwritten text datasets (the SCUT-HCCDoc and the SCUT-EPT) and one English handwritten text dataset (the IAM) show that the proposed method can achieve comparable results with the state-of-the-art methods. The code is available at https://github.com/Wukong90/3DAttention-MultiScaleTraining-for-HTR.
@@ 1 Introduction
There are three typical segmentation-free methods, i.e., hidden Markov model (HMM) [1, 2], connectionist temporal classification (CTC)[3, 4, 5, 6, 7, 8] and encoder–decoder (ED) framework [9, 10, 11, 12]. As shown in Fig. 1, in the HMM-based method, each character is modeled by an HMM and a text line can be represented by cascaded HMMs. A series of frames extracted from an original image by a left-to-right sliding window are assigned to the underlying states. Then, a neural network is used to estimate the posterior probabilities of the states, while the outputs of networks in CTC and ED-based approaches are character classes. In the CTC loss, a special character "blank" and a sophisticated rule are designed to split different characters, and the forward-backward algorithm can efficiently compute the probability of the underlying character sequence. In an encoder-decoder network, an image is usually fed into the encoder to generate the corresponding middle features. Based on the middle representations, the decoder is used to locate and predict the character sequence via attention mechanisms. The common cross-entropy loss can be directly used to adjust the parameters of the ED network.
@@
It is reasonable to explicitly model the 2D information of characters. However, it is very difficult to expand 1D HMM to 2D HMM [13] due to the computational complexity. Even for the CTC, ED-based approaches [14, 15, 6, 10], early networks just simply depend on the local receptive field of convolutional layers or recurrent units and gradually shrink the height of feature maps to small pixels via stacked pooling layers, which may lead to information loss. Although some attempts [16, 17, 11, 18] explicitly or implicitly employ 2D information in scene text recognition and character recognition, there is no obviously evidence shows that these method can be directly used in long handwritten text recognition. Obviously, the context information is important for the text recognition. In [19, 20, 21], the visual features and the linguistic knowledge are simultaneously integrated into a network. Actually, based on the visual features extracted from a certain image, the corresponding context information can also be naturally obtained by using recurrent units [5, 16, 22, 8] or transformers [17, 23, 24].
For the handwritten text recogniton task, recently, Hoang et al. [7] combined the radical-level CTC loss and the character-level loss while Ngo et al. [12] constructed a joint decoder to integrate the visual feature and the linguistic context feature. More recently, Peng et al. [25] built a full convolution network to simultaneously achieve the purpose of the character location, the character bounding boxes prediction and the character prediction. Lin et al. [26] propose a mobile text recognizer via searching the lightweight neural units. Furthermore, Peng et al. [27] detect and recognize characters in page-level handwritten text while Coquenet et al. [23] employ a fully convolutional network as the encoder with a transformer decoder for document recognition.
In this paper, inspired by the above three modeling ways, we propose a new recognition network by using a novel 3D attention module and global-local context information. The 3D attention module is employed to explicitly extract 2D information of block features with different resolutions. In detail, the 3D attention module is ingeniously decoupled into a 2D self-attention operation and a 1D attention-based aggregation operation. Based on the outputs (visual features) of the 3D attention module, we further extract the corresponding global-local context information via the self-attention and the recurrent unit, respectively. Finally, by integrating the visual features and the corresponding global-local context features, a well-designed representation for the recognition task can be obtained. In summary, the main contributions of this paper are as follows:
@@
    Main canonical neural units including attention mechanisms, fully-connected layer, recurrent unit and convolutional layer are ingeniously organized into a network.
@@
The remainder of this paper is organized as follows: Section 2 elaborates on the details of the proposed method. Section 3 reports the experimental results and analysis. Finally, we conclude the paper.
@@ 2 Methodology
As shown in Fig. 2, the proposed network includes three parts, i.e., the convolutional neural network (CNN), the 3D attention module and the features integration block. In the CNN, the hybrid attention module (HAM) [8] is used and each convolutional layer is equipped with the batch normalization [28]. For the outputs of the CNN, we employ a multi-scale framing strategy to extract different solutions. In this section, we elaborate on the details of the 3D attention module and the global-local context information. Moreover, the multi-scale training pipeline is also shown.
@@
${{\bf{p}}_{p,2i}}=\sin(p*{e^{2i*(-\ln(10000)/C)}})$ (3)
${{\bf{p}}_{p,2i+1}}=\cos(p*{e^{2i*(-\ln(10000)/C)}})$ (4)
where $i$ is indice along hidden dimensions. And then, we reshape each 3D block from ${C\times S\times H}$ to $SH\times C$ and conduct a self-attention operation for these sequential vectors ${\bf{f}}_{p}$ as Eq.5:
@@
2.2 Features Integration
Through the 3D attention module, a series of 3D feature maps can be transformed into the corresponding visual features ${\bf{r}}_{t}(t=0...T)$. As shown in Fig. 4-(a), for the visual feature ${\bf{r}}_{t}$, the global context feature ${\bf{l}}_{t}$ and the local context ${\bf{s}}_{t}$ can be extracted by using the self-attention mechanism and the recurrent units, respectively. In this paper, we directly employ the computation method of the ${\bf{f}}^{{}^{\prime}}$ in Fig. 3 to obtain the global context information ${\bf{l}}_{t}$. For the local context features, the standard LSTM (Fig. 4-(b)) is used. ${\bf{r}}_{t}$ represents the input at time $t$ and ${\bf{s}}_{t}$ is the corresponding output, $\bf{i}$, $\bf{g}$, $\bf{o}$ and $\bf{c}$ are the outputs of input gate, forget gate, output gate and cell vectors, respectively. Finally, the visual feature ${\bf{r}}_{t}$, the global feature ${\bf{l}}_{t}$ and the local feature ${\bf{s}}_{t}$ are simply contacted together:
@@ 2.2 Features Integration
Figure 4: The different features (visual features, global-local features) are combined before the classify layer and the details of the LSTM are illustrated.
@@ 2.3 Multi-Scale Training
The proposed multi-scale training include two parts. Firstly, as shown in Fig. 2, in the training stage, multiple feature sequences with different frame lengths are extracted. Although parallel branches including the same 3D attention block and global-local context block are used, we only retained the corresponding branch of the frame length 3 during the inference stage in our experiments.
Secondly, we employ the joint training of the CTC and the CE losses. Given the feature sequence $\bf{V}$ of a text line image, the text recognition task is to find the corresponding underlying n-character sequence $\bf{C}=\{C_{1},C_{2},...,C_{n}\}$, i.e, compute the posterior probability $p(\bf{C}|\bf{V})$. The CTC can be regarded as a loss function of neural networks:
@@
Figure 5: The decoder with attention. The Integration is similar to the module in Fig. 3
where $\bf{\Theta}$ represents the parameters of the recognition network and $\bm{\pi}$ is the predicted character sequence under the constraint of $\varphi({\bm{\pi}})={\bf{C}}$. The function $\bm{\pi}$ only retains one of the consecutive adjacent repeating characters and removes the ‘blank’ characters. The CTC loss (Eq.10) can be efficiently computed by using the forward-backward algorithm [3]. For the auxiliary decoder in Fig. 5, the CE loss directly predicts the probability $p(C_{t}|{\bf{x}}_{t})$:
@@
0:    The randomly initialized parameter sets $\bf{\Theta}$, $\bf{\Gamma}$}; The loss functions $L_{\rm{CTC}}$, $L_{\rm{CE}}$ and the coressponding weight coefficients ${\lambda_{1},\lambda_{2}}$ in the training stage of the main recognition network (MRN($\bf{\Theta}$)), and the auxiliary decoder (Dec($\bf{\Gamma}$)), respectively.
@@
2:  Jointly train the parameter set ${\bf{\Theta}},{\bf{\Gamma}}$ based on the CTC and the CE losses.$L={\lambda_{1}}L_{\rm{CTC}}+{\lambda_{2}}L_{\rm{CE}}$ $\bf{\Theta,\Gamma}=\rm{Adam}({\bf{\Theta,\Gamma}},L)$
3:  return  The MRN($\bf{\Theta}$).
@@ 3.1 Datasets and Evaluation Metrics
Table 2: The standard partition of the dataset IAM.
@@
For the English handwritten text recognition, we evaluate the performance of our method on the IAM dataset. The IAM dataset contains a total number of 9,862 text lines written by 500 writers. It provides one training set, one test set and two validation sets. The text lines of all data sets are mutually exclusive, thus, each writer has contributed to one set only.
Table 1 lists the number of text lines, characters and classes in both Chinese datasets. In the SCUT-EPT dataset, we removed 681 training text images including abnormal structures, i.e. characters swapping and overlap. Similarly, in the SCUT-HCCDoc dataset, some low-quality images were removed and only 91,261 text instances were used in the training set. All test images in both datasets were evaluated. Table 2 lists the standard partition of the English dataset  [^1] https://fki.tic.heia-fr.ch/databases/iam-handwriting-database. All training images, validation images and test images were used in our experiments and the best model was selected for testing according to the results on the validation sets. The character accuracy rate (AR) was used to evaluate the performance of the text recognition model. The evaluation criterion is defined as follows:
@@ 3.2.1 Ablation Experiments
We only compare the results of different frame lengths and all sliding windows have the same height. Table 3 shows the different results of different frame resolutions. We can observe that frame length 3 is the optimal configuration for all experiments, which means the corresponding window size is suitable for most characters. For a longer or shorter feature sequence, it is difficult to discriminate different regions of characters. For example, when we used a small frame length, the Chinese recognition results obviously decreased from 88.8% to 88.06%, 76.67% to 75.13%, respectively, while the English recognition performance only changed from 94.17% to 94.11%, which may be owing to a smaller width for most English characters. Furthermore, we can obtain consistently improvements by integrating multiple resolutions, i.e., the ARs achieve 89.16%, 76.96% and 94.37%, respectively. As shown in Fig. 2, in the training stage, although three 3D attention blocks and global-local context blocks were used in parallel, we only retained the corresponding branch of the frame length 3 during the inference stage.
@@
Based on the optimal frame length and single-scale training, Table 4 lists the results with/without the 3D attention module and the global-local context information. If we did not adopt the 3D attention module in experiments, a pooling operation was directly employed to reduce the height of the last convolutional feature maps to 1. The 3D attention module can steadily improve the Chinese recognition performances. On the SCUT-HCCDoc data set, the AR is improved from 88.51% to 88.8%. On the SCUT-EPT data set, the AR increases from 75.00% to 76.67%. Although it seems that the English recognition without the 3D attention module has a higher AR (94.31% vs. 94.17%), the network based on the multi-resolutions training still obtains the best AR. As the conclusions demonstrated by many previous researches, there exists an obvious gap with/without using the context information for the sequence task. The results dramatically decrease from 88.8% to 85.36%, 76.67% to 72.28% and 94.17% to 87.0%, respectively. It is interesting to observe that the pure CNN-based models without the 3D attention block and the context information can also achieve 85.76%, 72.53% and 85.93%, respectively.
@@
$\checkmark$ $\checkmark$ $\bm{88.8}$ $\bm{76.67}$ 94.17%
$\checkmark$ $\bm{\times}$ 85.36 72.28 87.0%
$\bm{\times}$checkmark 88.51 75.00 $\bm{94.31\%}$
@@
In Table 5, we compare the results of different training strategies, i.e. the weights are only optimized by the CTC loss or jointly trained by the CTC loss and the CE loss. Although the Chinese decoding results of the auxiliary network are not good, the main recognition network still benefits from the joint training. The best network can achieve 89.34%, 77.15% and 94.41%, respectively. Moreover, we also list the storage and the running time comparisons of different methods. In order to make a fair comparison, all experiments were evaluated on the same machine and the experimental configurations on the same dataset were consistent. The decoding time of the CTC is defined as 1. Generally speaking, the results show that the ED method needs more time consumption and storage.
@@
Table 6: Performance comparison of our proposed method and other state-of-the-arts methods. For the English handwritten recognition task, all listed results do not use any lexicons.

### v3

@@ Abstract
The segmentation-free research efforts for addressing handwritten text recognition can be divided into three categories: connectionist temporal classification (CTC), hidden Markov model and encoder–decoder methods. In this paper, inspired by the above three modeling methods, we propose a new recognition network by using a novel three-dimensional (3D) attention module and global-local context information. Based on the feature maps of the last convolutional layer, a series of 3D blocks with different resolutions are split. Then, these 3D blocks are fed into the 3D attention module to generate sequential visual features. Finally, by fusing the visual features and the corresponding global-local context features, a well-designed representation can be obtained. Main canonical neural units including attention mechanisms, fully-connected layers, recurrent units and convolutional layers are efficiently organized into a network and can be jointly trained by the CTC loss and the cross-entropy loss. Experiments on the latest Chinese handwritten text datasets (the SCUT-HCCDoc and the SCUT-EPT) and one English handwritten text dataset (the IAM) show that the proposed method can achieve comparable results with the state-of-the-art methods. The code is available at https://github.com/Wukong90/3DAttention-MultiScaleTraining-for-HTR.
@@ 1 Introduction
In text recognition tasks, there are three typical segmentation-free methods, i.e., hidden Markov model (HMM) [1, 2], connectionist temporal classification (CTC)[3, 4, 5, 6, 7, 8] and encoder–decoder (ED) framework [9, 10, 11, 12]. As shown in Fig. 1, in the HMM-based method, each character is modeled by an HMM and a text line can be represented by cascaded HMMs. A series of frames extracted from an original image by a left-to-right sliding window are assigned to the underlying states. Then, a neural network is used to estimate the posterior probabilities of the states, while the outputs of networks in CTC and ED-based approaches are character classes. In the CTC loss, a special character "blank" and a sophisticated rule are designed to split different characters, and the forward-backward algorithm can efficiently compute the probability of the underlying character sequence. In an encoder-decoder network, an image is usually fed into the encoder to generate the corresponding middle features. Based on the middle representations, the decoder is used to locate and predict the character sequence via attention mechanisms. The common cross-entropy loss can be directly used to adjust the parameters of the ED network.
@@
It is reasonable to explicitly model the 2D information of characters. However, it is very difficult to expand 1D HMM to 2D HMM [13] due to the computational complexity. Even for the CTC, ED-based approaches [14, 15, 6, 10], early networks usually simply depend on the local receptive field of convolutional layers or recurrent units and gradually shrink the height of feature maps to small pixels via stacked pooling layers, which may lead to information loss. Although some attempts [16, 17, 11, 18] explicitly or implicitly use 2D information in scene text recognition and character recognition, there is no obvious evidence showing that these methods can be directly used in long handwritten text recognition. Obviously, the context information is important for the text recognition. In [19, 20, 21], the visual features and the linguistic knowledge are simultaneously integrated into a network. Actually, based on the visual features extracted from a certain image, the corresponding context information can also be naturally obtained by using recurrent units [5, 16, 22, 8] or transformers [17, 23, 24].
For the handwritten text recognition task, recently, Hoang et al. [7] combined the radical-level CTC loss and the character-level loss while Ngo et al. [12] constructed a joint decoder to integrate the visual feature and the linguistic context feature. More recently, Peng et al. [25] built a full convolution network to simultaneously achieve the purpose of the character location, the character bounding boxes prediction and the character prediction. Lin et al. [26] propose a mobile text recognizer via searching the lightweight neural units. Furthermore, Peng et al. [27] detect and recognize characters in page-level handwritten text while Coquenet et al. [23] use a fully convolutional network as the encoder with a transformer decoder for document recognition.
In this paper, inspired by the above three modeling ways, we propose a new recognition network by using a novel 3D attention module and global-local context information. The 3D attention module is employed to explicitly extract 2D information of block features with different resolutions. In detail, the 3D attention module is ingeniously decoupled into a 2D self-attention operation and a 1D attention-based aggregation operation. Based on the outputs (visual features) of the 3D attention module, we further extract the corresponding global-local context information via the self-attention and the recurrent unit, respectively. Finally, by fusing the visual features and the corresponding global-local context features, a well-designed representation for the recognition task can be obtained. In summary, the main contributions of this paper are as follows:
@@
    Main canonical neural units including attention mechanisms, fully-connected layers, recurrent units and convolutional layers are ingeniously organized into a network.
@@
The remainder of this paper is organized as follows: Section 2 elaborates on the details of the proposed method. Section 3 reports the experimental results and analysis. Finally, we discuss the advantages of the proposed method and conclude the paper.
@@ 2 Methodology
As shown in Fig. 2, the proposed network includes three parts, i.e., the convolutional neural network (CNN), the 3D attention module and the features fusion block. In the CNN, the hybrid attention module (HAM) [8] is used and each convolutional layer is equipped with the batch normalization [28]. For the outputs of the CNN, we use a multi-scale framing strategy to extract different solutions. In this section, we elaborate on the details of the 3D attention module and the global-local context information. Moreover, the multi-scale training pipeline is also shown.
@@
${{\bf{p}}_{p,2i}}=\sin(p{e^{2i(-\ln(10000)/C)}})$ (3)
${{\bf{p}}_{p,2i+1}}=\cos(p{e^{2i(-\ln(10000)/C)}})$ (4)
here $i$ is the index along hidden dimensions. And then, we reshape each 3D block from ${C\times S\times H}$ to $SH\times C$ and conduct a self-attention operation for these sequential vectors ${\bf{f}}_{p}$ as Eq.5:
@@
2.2 Features Fusion
Through the 3D attention module, a series of 3D feature maps can be transformed into the corresponding visual features ${\bf{r}}_{t}(t=0...T)$. As shown in Fig. 4-(a), for the visual feature ${\bf{r}}_{t}$, the global context feature ${\bf{l}}_{t}$ and the local context ${\bf{s}}_{t}$ can be extracted by using the self-attention mechanism and the recurrent units, respectively. We directly use the computation method of the ${\bf{f}}^{{}^{\prime}}$ in Fig. 3 to obtain the global context information ${\bf{l}}_{t}$. For the local context features, the standard LSTM (Fig. 4-(b)) is used. ${\bf{r}}_{t}$ represents the input at time $t$ and ${\bf{s}}_{t}$ is the corresponding output, $\bf{i}$, $\bf{g}$, $\bf{o}$ and $\bf{c}$ are the outputs of input gate, forget gate, output gate and cell vectors, respectively. Finally, the visual feature ${\bf{r}}_{t}$, the global feature ${\bf{l}}_{t}$ and the local feature ${\bf{s}}_{t}$ are simply contacted together:
@@ 2.2 Features Fusion
Figure 4: The different features (visual features, global-local features) are combined before the classification layer and the details of the LSTM are illustrated.
@@ 2.3 Multi-Scale Training
The proposed multi-scale training includes two parts. Firstly, as shown in Fig. 2, in the training stage, multiple feature sequences with different frame lengths are extracted. Although parallel branches including the same 3D attention block and global-local context block are used, we only retained the corresponding branch of the frame length 3 during the inference stage in our experiments.
Secondly, we use the joint training of the CTC and the CE losses. Given the feature sequence $\bf{V}$ of a text line image, the text recognition task is to find the corresponding underlying n-character sequence ${\bf{C}}=\{C_{1},C_{2},...,C_{n}\}$, i.e, compute the posterior probability $p(\bf{C}|\bf{V})$. The CTC can be regarded as a loss function of neural networks:
@@
Figure 5: The decoder with attention. The Attention Weight Computation module is similar to the Integration module in Fig. 3
where $\bf{\Theta}$ represents the parameters of the recognition network and $\bm{\pi}$ is the predicted character sequence under the constraint of $\varphi({\bm{\pi}})={\bf{C}}$. The function $\varphi$ only retains one of the consecutive adjacent repeating characters and removes the ‘blank’ characters. The CTC loss (Eq.10) can be efficiently computed by using the forward-backward algorithm [3]. For the auxiliary decoder in Fig. 5, the CE loss directly predicts the probability $p(C_{t}|{\bf{x}}_{t})$:
@@
[…]
@@ 3.1 Datasets and Evaluation Metrics
Table 2: The standard partition of the IAM dataset.
@@
For the English handwritten text recognition, we evaluate the performance of our method on the IAM dataset. The IAM dataset contains a total of 9,862 text lines written by 500 writers. It provides one training set, one test set and two validation sets. The text lines of all data sets are mutually exclusive, thus, each writer has contributed to one set only.
Figure 6: Some removed images in the SCUT-HCCDoc dataset are shown above the dashed line. They are very low-quality (illegible text, random graffiti, high overlap of text and background, and incomplete characters) or vertical writing text. The lower part of the dashed line shows typical removed images in the SCUT-EPT dataset. These images include abnormal structures, i.e., characters swapping and overlap, and more details can be found in [34].
Table 1 lists the number of text lines, characters and classes in both Chinese datasets. In the SCUT-EPT dataset  [^1] https://github.com/HCIILAB/SCUT-EPT_Dataset_Release. Prof. Lianwen Jin (eelwjin@scut.edu.cn) at South China University of Technology (SCUT) is the owner of this dataset., we removed 681 training text images including abnormal structures, i.e., characters swapping and overlap. Similarly, in the SCUT-HCCDoc dataset  [^2] https://github.com/HCIILAB/SCUT-HCCDoc_Dataset_Release. Prof. Lianwen Jin (eelwjin@scut.edu.cn) at SCUT is the owner of this dataset. Thanks to his team for building two useful datasets., some very low-quality (illegible text, random graffiti, high overlap of text and background, and incomplete characters) or vertical writing text images were removed and only 91,261 text instances were used in the training set. All test images in both datasets were evaluated. Table 2 lists the standard partition of the English dataset  [^3] https://fki.tic.heia-fr.ch/databases/iam-handwriting-database. Accessed 2021.. All training images, validation images and test images were used in our experiments and the best model was selected for testing according to the results on the validation sets. The character accuracy rate (AR) was used to evaluate the performance of the text recognition model. The evaluation criterion is defined as follows:
@@ 3.2.1 Ablation Experiments
We only compare the results of different frame lengths and all sliding windows have the same height. Table 3 shows the different results of different frame resolutions. We can observe that frame length 3 is the optimal configuration for all experiments, which means the corresponding window size is suitable for most characters. For a longer or shorter feature sequence, it is difficult to discriminate different regions of characters. For example, when we used a small frame length, the Chinese recognition results obviously decreased from 88.8% to 88.06%, 76.67% to 75.13%, respectively, while the English recognition performance only changed from 94.17% to 94.11%, which may be owing to a smaller width for most English characters. Furthermore, we can obtain consistent improvements by integrating multiple resolutions, i.e., the ARs achieve 89.16%, 76.96% and 94.37%, respectively. As shown in Fig. 2, in the training stage, although three 3D attention blocks and global-local context blocks were used in parallel, we only retained the corresponding branch of the frame length 3 during the inference stage.
@@
Based on the optimal frame length and single-scale training, Table 4 lists the results with/without the 3D attention module and the global-local context information. If we did not adopt the 3D attention module in experiments, a pooling operation was directly used to reduce the height of the last convolutional feature maps to 1. The 3D attention module can steadily improve the Chinese recognition performance. On the SCUT-HCCDoc data set, the AR is improved from 88.51% to 88.8%. On the SCUT-EPT data set, the AR increases from 75.00% to 76.67%. Although it seems that the English recognition without the 3D attention module has a higher AR (94.31% vs. 94.17%), the network based on the multiple-resolutions training still obtains the best AR. As the conclusions demonstrated by many previous researches, there exists an obvious gap with/without using the context information for the sequence task. The results dramatically decrease from 88.8% to 85.36%, 76.67% to 72.28% and 94.17% to 87.0%, respectively. It is interesting to observe that the pure CNN-based models without the 3D attention block and the context information can also achieve 85.76%, 72.53% and 85.93%, respectively.
@@
$\checkmark$ $\checkmark$ $\bm{88.80}$ $\bm{76.67}$ 94.17
$\checkmark$ $\bm{\times}$ 85.36 72.28 87.00
$\bm{\times}$checkmark 88.51 75.00 $\bm{94.31}$
@@
In Table 5, we compare the results of different training strategies, i.e., the weights are only optimized by the CTC loss or jointly trained by the CTC loss and the CE loss. Although the Chinese decoding results of the auxiliary network are not good, the main recognition network still benefits from the joint training. The best network can achieve 89.34%, 77.15% and 94.41%, respectively. Moreover, we also list the storage and the running time comparisons of different methods. In order to make a fair comparison, all experiments were evaluated on the same machine and the experimental configurations on the same dataset were consistent. In Table 5, the decoding time of the CTC is defined as 1. Generally speaking, considering the auxiliary decoder with attention needed and the cyclic decoding way used during the inference stage, it is reasonable that the ED method needs more time consumption and storage. As shown in Fig 2, for the joint training strategy of CTC and CE, only the faster CTC method is used in the test stage, i.e., it has the same time consumption with the pure CTC training strategy.
@@
Table 6: Performance comparison of our proposed method and other state-of-the-art methods. For the English handwritten recognition task, all listed results do not use any lexicons.